OpenStudy (anonymous):

METAL WILL BE REWARDED Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1))

OpenStudy (anonymous):

g(-1)=1+7-9=-1 f(-1)=-1-2=-3